# AI, Climate, and Regulation: From Data Centers to the AI Act

### v1

@@ Abstract
We live in a world that is experiencing an unprecedented boom of AI applications that increasingly penetrate and enhance all sectors of private and public life, from education, media, medicine, and mobility to the industrial and professional workspace, and—potentially particularly consequentially—robotics. As this world is simultaneously grappling with climate change, the climate and environmental implications of the development and use of AI have become an important subject of public and academic debate. In this paper, we aim to provide guidance on the climate-related regulation for data centers and AI specifically, and discuss how to operationalize these requirements. We also highlight challenges and room for improvement, and make a number of policy proposals to this end. In particular, we propose a specific interpretation of the AI Act to bring reporting on the previously unadressed energy consumption from AI inferences back into the scope. We also find that the AI Act fails to address indirect greenhouse gas emissions from AI applications. Furthermore, for the purpose of energy consumption reporting, we compare levels of measurement within data centers and recommend measurement at the cumulative server level. We also argue for an interpretation of the AI Act that includes environmental concerns in the mandatory risk assessment (sustainability risk assessment, SIA), and provide guidance on its operationalization. The EU data center regulation proves to be a good first step but requires further development by including binding renewable energy and efficiency targets for data centers. Overall, we make twelve concrete policy proposals, in four main areas: Energy and Environmental Reporting Obligations; Legal and Regulatory Clarifications; Transparency and Accountability Mechanisms; and Future Far-Reaching Measures beyond Transparency. [^1] Date of this version: Oct 8, 2024.
1 Introduction
The environmental consequences of artificial intelligence (AI) are becoming ever more apparent as large models are increasingly trained and deployed across society. This inherently intertwines the digital transformation with questions of climate change, particularly concerning the energy and water consumption of AI models. These features are attracting attention from both the public and academia [11, 10, 3, 7, 18, 4, 13]. Indeed, as AI is integrated into ever more processes—from economic operations, logistics and employment to administrative and judicial proceedings—agentic systems are being deployed, and machine learning is increasingly writing code for robots and steering them in real time, limiting the climate effects of AI throughout the value chain is becoming an ever more pressing societal and regulatory concern.
Starting at the hardware side of the AI value chain, one major, broadly debated component is energy efficiency targets for data centers, [^2] Further EU legislation is anticipated, following the Commission’s report by 15 May 2025 under Art. 12(5) of the Energy Efficiency Directive (EU) 2023/1791. with concerns growing that the supply of renewable energy may not keep up with their increasing demand [1]. Leading companies like Google have already sounded the alarm, noting that rising energy demands from AI may jeopardize their carbon-zero commitments. [^3] Google 2024: Environmental Report.
Regulation is likely to play a pivotal role in addressing the collective action problem posed by AI’s impact on the climate. Globally, various initiatives are under development to create comprehensive legal frameworks for AI. One of the most notable examples is the EU AI Act, which took effect in early August 2024. This legislation includes substantial provisions aimed at climate considerations, with specific requirements for reporting, risk assessments, and mitigation efforts. Consequently, it is anticipated that AI’s environmental footprint will become a critical market factor, influence reputational outcomes, and enable enhanced public scrutiny by stakeholders such as NGOs and analysts. Furthermore, the revised Energy Efficiency Directive, along with the Delegated Regulation, introduces regulatory measures targeting data center operators.
This paper explores the current and proposed regulatory landscape governing data center operators, along with the AI Act’s climate-related provisions, from both legal and technical perspectives. Our analysis focuses on three key contributions. First, we identify significant gaps in the AI Act and, to some extent, in the Energy Efficiency Directive. Second, we highlight the operational difficulties in implementing the AI Act’s current mandates. Finally, we offer a set of policy recommendations aimed at overcoming these challenges. These reforms extend beyond the scope of the AI Act, referencing broader initiatives such as the recently adopted Global Digital Compact at the United Nations.
2 Technical Background
To lay the ground for the subsequent legal and technical analysis, we start by introducing some core technical aspects of AI models.
@@ 2 Technical Background
From a technical perspective, it is important to distinguish between training, fine-tuning, and inference. Training, on the one hand, refers to the process of initially adjusting a model’s parameters or weights to fit the data. This process is highly compute-intensive and typically requires a significant amount of energy [12, 11, 21, 7]. The terms training and pre-training are used interchangeably. The pre-training step involves training an LLM on large-scale data. LLM performance strongly depends on the model scale (number of parameters), which also requires more training data [9, 17]. Therefore, a continuous scaling of the models and thus also of the energy and resource consumption is to be expected, which can also be observed through the continuous publication of new and ever larger models. It is currently not possible to assess when an actual saturation of model performance will occur [19], as evaluation metrics, available data and tasks are also subject to ongoing adaptation and development.
@@
Fine-tuning aims at adjusting a pretrained model to fit more specific data in order to perform better at specific tasks. It is usually performed by retraining specific parts or even the full model. One can also add layers or model architecture components to better account for the new data or task. Overall, fine-tuning follows pre-training and involves a significantly smaller amount of training data and compute budget. The specific procedure for successful fine-tuning is very model-, data-, and task-specific, and rather empirical. Therefore, the computational and energy-intensity can vary significantly for this step. For example, as highlighted by Luccioni et al. [12], the energy usage for fine-tuning the Bloomz-7B required 7,571 kWh compared to 51,686 kWh for the entire training process, adding another 15 % to the inital consumption. Fine-tuning and other related methods enable researchers and companies to make use of large-scale models which would otherwise be prohibitively expensive.
Inference, on the other hand, refers to the procedure of generating a prediction or output from a trained model based on provided input features or data. In practical applications, models generate inferences during deployment. For example, prompting an LLM constitutes an inference. Typically, an individual inference consumes little energy compared to training; but there are many more inference than training events.
From a hardware perspective, the training and fine-tuning of large models generally requires a distributed arrangement of many servers, each handling subsets of the model and data. Techniques are used to synchronize these distributed components. Presently, the hardware predominantly consists of Nvidia graphics processing unit (GPU)-intensive servers.
Inference in large models is also conducted using GPU-based servers. Although certain models can be accommodated within the memory of a single GPU, inferences for most large models surpass the GPU memory capacity and must also be executed in a distributed manner. It is important to mention that methods like quantization can be employed to lower the computational and memory requirements during inference. This technique involves using model weights at a reduced precision. Also, task-specific models tend to consume less energy than general-purpose models [12]. This demonstrates that energy and hardware requirements are not fixed for the very same task. Given the same performance necessities, more energy-efficient and hardware-efficient models might be, and increasingly become, available.
Research on efficient LLM training and fine-tuning is a growing field with a large variety of approaches involving the full lifecycle of a model. Examples include mixed precision pre-training [8, 17], efficient resource scheduling utilizing thousands of GPUs [15], low-rank adaptation [6], among others [20]. Nevertheless, this research concentrates on computing budget rather than energy efficiency, despite a significant relationship between them.
Energy usage in hardware can be classified into static and dynamic energy [16, 2]. Static energy is the power required to run hardware devices without any specific load (not limited to GPUs). Dynamic energy is consumed when transistors switch their binary states and are electrically charged, driving the predominant portion of power consumption. Typically, fully utilizing hardware devices is more energy-efficient than operating them at lower workloads for extended periods since static energy consumption remains independent of workload.
@@
Most of these computations are conducted in data centers. The power usage effectiveness (PUE) metric reflects the energy efficiency of a data center. It indicates the ratio of the total energy needed by a data center, including components like cooling, to the energy used solely by the computational devices. A PUE of 1.0 would imply ideal efficiency, meaning the data center uses only the energy necessary to power the computational devices. Conversely, a PUE of 1.2 indicates that in order to use 1kWh for computing hardware, the total energy consumption of the data center is 1.2kWh. The average data center PUE in 2023 was 1.58 globally [^4] Statista and 1.6 in the EU [^5] COM Joint Research Centre.
3 Regulation on Data Centers
@@
Data centers run all kinds of operations, such as cloud computing, crypto currencies and the Internet at large. Recently, AI training and inference have experienced massive growth. Broad data center regulation, therefore, indirectly governs environmental effects of AI and constitutes the backdrop against which specific AI regulation must be viewed. Taking a look at the requirements for data centers can help to find a coherent and effective interpretation of the requirements in the AI Act as it builds on already available data and established methodology.
EU rules stipulate data collection and reporting obligations for data centers (see subsection 3.1) and mandatory energy audits (subsection 3.2). However, they lack binding efficiency and renewable energy targets. Most of these EU rules need national implementation by Member States. Notably, the German implementation goes much further by broadening the scope and putting in place additional requirements, including energy efficiency and renewable energy targets for data centers (subsection 3.3).
In the U.S., a bill for an AI Environmental Impacts Act was introduced in the Senate, but has not yet been encacted (subsection 3.4). The proposed law takes a very soft approach. It does not suggest any binding rules for the industry, but would commission Federal Agencies to prepare the ground for future industry legislation. Any reporting under the law would be voluntary.
3.1 EU Data Collection and Reporting Obligations for Data Centers
@@
In the EU, data collection and reporting obligations for data centers were established by two recent legal acts, Art. 12 of the recast Energy Efficiency Directive EU/2023/1791 of September 13, 2023 ("EED"), and the Commission Delegated Regulation EU/2024/1364 of March 14, 2024 ("Delegated Regulation"). The new rules apply to all data centers in the EU with a power demand of the installed information technology (IT) of at least 500kW, which includes small-sized data centers. Data center operators are required to collect, make publicly available and report to a EU database information that is deemed relevant for the sustainability assessment of the data centers and the industry as a whole. The reporting is mandated on an annual basis.
@@
The required data includes energy consumption, power utilization, temperature set points, waste heat utilisation, water usage and use of renewable energy (EED, Annex VII(c)). Notably, while the EED focuses on energy and power, the reporting of water usage is a significant step forward as both energy and water consumption have raised concerns in AI settings [10, 12]. In addition, the Water Framework Directive can be harnessed to limit the overall amount water a data center may consume, and also control for any potential loss of water quality [4].
@@
The Delegated Regulation provides specific key performance indicators and methodology. Most notable is the requirement to measure and report the energy consumption of the installed information technology. Following the standard-methodology for the calculation of PUE, [^6] The Delegated Regulation refers to the European standard CEN/CENELEC EN 50600-4-2. the energy consumption must be measured at the uninterruptible power system (UPS) or, if not existent, at the power distribution unit (PDU) or at another point specified by the data center (see Delegated Regulation, Annex II(1)(e)). These options are depicted in Figure 1 as Categories 1-3.
Figure 1: Measuring points for IT equipment energy consumption pursuant to the Delegated Regulation, Annex II
Prior to the issuance of the Delegated Regulation, the Commission conducted an expert study on the options for a reporting scheme, which the Commission broadly followed. [^7] European Commission, Reporting requirements on the energy performance and sustainability of data centers for the Energy Efficiency Directive, Task A report: Options for a reporting scheme for data centers. Even though it might be more accurate to measure the energy consumption closer to the ICT equipment, the study recommends the UPS as a primary measuring point. [^8] Ibid, p. 24-25. Measuring directly at the ICT equipment is deemed impractical because colocation data centers have no direct access to the ICT equipment housed for their customers. Measuring at PDU level is also considered impractical because many data centers do not yet have the equipment to measure energy consumption after the UPS. The UPS is also the easiest point to implement measuring for those who have no measurement in place yet, at all. Overall, it is deemed good enough to identify data centers with a high energy reduction potential of the supporting infrastructure. The three-tiered approach therefore seems like a pragmatic compromise, taking into account the technical constraints of colocation data center operators, in particular. Measuring at the UPS, or at the PDU alternatively, is not expected to make a significant difference because, typically, no additional energy consuming equipment is interlaced between those two measuring points.
@@
While the EED sets reduction targets for the Union and Member State economies as a whole, it lacks efficiency and renewable energy targets specifically for data centers. Pursuant to Art. 12(5) EED the Commission must report to the European Parliament and Council by May 15, 2025 an assessment of the hitherto reported data and legislative proposals for further measures. Proposals are to establish minimum performance standards, including a timeframe, and an assessment of the feasability of a transition to net-zero emissisons in the data center sector.
3.2 Energy Management Systems and Energy Audits
@@
The EED also requires that Member States mandate companies with an average annual energy consumption of more than 10 TJ to conduct an energy audit at least every four years and those with a consumption of more than 85 TJ to implement an energy management system including regular energy audits (Art. 11 EED). This would also apply to operators of data centers. The Directive sets up certain minimum criteria for energy audits (Annex VI EED) and refers to the relevant international or European standards (Recital 80 EED). The legal minimum criteria, however, do not dictate methodology on measurement of energy consumption.
3.3 German 2023 Energy Efficiency Act
In Germany, the Energy Efficiency Act of 8 Nov 2023 implements the EED and establishes a national reporting scheme and additional requirements, including specific efficiency and renewable energy targets for data centers.

### v2

@@ Abstract.
We live in a world that is experiencing an unprecedented boom of AI applications that increasingly penetrate all sectors of private and public life, from education, media, medicine, and mobility to the industrial and professional workspace. As this world is simultaneously grappling with climate change, the climate effects and environmental implications of the development and use of AI have become an important subject of public and academic debate. In this paper, we aim to provide guidance on the climate-related regulation for data centers and AI specifically, and discuss how to operationalize these requirements. We also highlight challenges and room for improvement, and make a number of policy proposals to this end. In particular, we propose a specific interpretation of the AI Act to bring reporting on the previously unaddressed energy consumption from AI inferences back into the scope. We also find that the AI Act fails to address indirect greenhouse gas emissions from AI applications. Furthermore, for the purpose of energy consumption reporting, we compare levels of measurement within data centers and recommend measurement at the cumulative server level. We also argue for an interpretation of the AI Act that includes environmental concerns in the mandatory risk assessments (sustainability risk assessment, SIA), and provide guidance on its operationalization. The EU data center regulation proves to be a good first step but requires further development by including binding renewable energy and efficiency targets for data centers. Overall, we make twelve concrete policy proposals, in four main areas: Energy and Environmental Reporting Obligations; Legal and Regulatory Clarifications; Transparency and Accountability Mechanisms; and Future Far-Reaching Measures beyond Transparency.
1. Introduction
The environmental consequences of artificial intelligence (AI) are becoming ever more apparent as large models are increasingly trained and deployed across society. This inherently intertwines the digital transformation with questions of climate change, particularly concerning the energy and water consumption of AI models. These features are attracting attention from both the public and academia (see below, 2.). Concerns are growing that the supply of renewable energy may not keep up with their increasing demand triggered by the AI scaling race (Anonymous, 2024a) in which the EU is poised to join the U.S. and China under the EU Commission’s new AI Continent Action Plan (European Commission, 2025). Scholars and activists, in turn, are increasingly critiquing and situating digital technology within the framework of environmental justice (Kazansky et al., 2022; Sotolongo, 2023; Yang et al., 2025; Rakova and Dobbe, 2023; Foundation, 2024). Work in political economy has shed a light on the extractive nature of AI-based socio-technical systems, extracting and, at times, exploiting, both natural and human resources (Keskin and Kiggins, 2021; Kasy, 2024; Ekbia and Nardi, 2017; Crawford, 2021; Mills, 2024), often particularly from marginalized communities (Perrigo, 2023; Noble, 2018) and regions (Li et al., 2023c; Yang, 2021). Both the quest for performance and the funding logic behind AI force providers to scale them to a point where this trajectory may push against planetary limits (Brevini, 2021; Crawford, 2021; Brevini, 2024; Bender et al., 2021).
Regulatory frameworks are beginning to address these challenges, too. While the new U.S. administration scraps environmental rules and generally deregulates, the EU Artificial Intelligence Act (AI Act) becomes applicable as the world’s most comprehensive attempt at direct AI regulation. It features environmental protection as one of its core goals and contains dedicated sustainability rules, which also apply to US and other non-EU providers offering models in the EU. Similarly, the Digital Services Act (DSA) compels Very Large Online Platforms and Very Large Online Search Engines to thoroughly assess and mitigate systemic risks, which is of particular relevance for hybrid platforms increasingly integrating AI. As scholars have pointed out, this includes climate and sustainability risks (Buri and Griffin, 2024), and intersects partially with obligations under the AI Act (Hacker, 2024a; Helberger and Diakopoulos, 2023). In parallel, the Energy Efficiency Directive (EU) 2023/1791 introduces transparency obligations for data centers, a critical infrastructure supporting AI. However, these regulatory instruments were developed independently, leading to fragmentation and gaps. The AI Act originally pursued climate-related objectives especially in the European Parliament’s position of June 2023, but over successive drafts and the trilogue negotiations, its environmental provisions were diluted, particularly with regard to energy consumption from AI inference and risk assessments. Meanwhile, the EU’s data center regulation remains incomplete, lacking binding efficiency and renewable energy targets. These shortcomings threaten to undermine the most promising legislative avenues for tackling AI’s climate effects. The EU’s recent push for massive investments in computing infrastructure under the AI Continent Action Plan, including the establishment of AI Gigafactories (European Commission, 2025), is a late but necessary impetus for technological and strategic autonomy but must not endanger the Union’s climate goals under the Green New Deal.
Against this background, this paper makes three key contributions. First, we offer the first thorough analysis of the AI Act’s final version from an environmental perspective. Unlike prior scholarship, which has examined these areas separately, we also scrutinize the interplay between data center regulation and direct AI regulation. AI sustainability is not governed by a single legal instrument but rather emerges from the interaction of multiple frameworks and actors. This fragmented approach creates legal uncertainty and regulatory loopholes — notably the omission of AI inference from reporting obligations and the failure to mandate Sustainability Impact Assessments (SIA). We then propose a novel interpretation of the AI Act to include AI inferences in the reporting obligations of general-purpose AI providers, and to incorporate sustainability considerations into the mandatory risk assessment framework to effectuate the goal of environmental protection enshrined in Art. 1 and Recitals 1, 2 and 8.
Second, we examine the practical challenges of implementing AI-related environmental obligations and provide technical recommendations. The crux here is to operationalize existing rules and make climate reporting work by developing metrics that are both meaningful and workable in practice.
Third, we identify areas where the AI Act and related regulatory frameworks require future amendments. In light of upcoming evaluation rounds, we advocate for closing loopholes and reinstating environmental obligations that were weakened during the legislative process. While the current political momentum seems to favor deregulation, the climate emergency persists. Therefore, it is crucial to improve the best legal tools still available, and to advocate for a nuanced legal framework to bring the twin transitions of AI and climate together. Our twelve policy proposals address four key areas: Energy and Environmental Reporting Obligations; Legal and Regulatory Clarifications; Transparency and Accountability Mechanisms; and Future Far-Reaching Measures beyond Transparency.
2. Related Work
Due to the interdisciplinary nature of sustainable AI, related works originate from three main fields—computer science; humanities and policy; and legal research. In recent years, the computer science and AI community has become increasingly aware of the environmental effects of AI, which led to both policy-oriented (Schwartz et al., 2020; Freitag et al., 2021; Chien, 2021; Bender et al., 2021; Knowles, 2021; Taddeo et al., 2021; Kaack et al., 2022; Knowles et al., 2022; Luccioni et al., 2024c, a, 2025) and technical contributions, particularly on the environmental effects of data centers operations and AI-related computing (OECD, 2022; Guidi et al., 2024; Aljbour et al., 2024), studies for tracking (Dayarathna et al., 2015; Luccioni et al., 2024b; Alder and Herbrich, 2024; Luccioni and Hernandez-Garcia, 2023; Luccioni et al., 2023; Wu et al., 2022; Dodge et al., 2022; Patterson et al., 2021; Zhou et al., 2020) and techniques for reducing the emissions of data centers and AI models (Beloglazov et al., 2011; Koomey and Masanet, 2021; Güler and Yener, 2021; Ludvigsen, 2022; Kneese and Young, 2024). The water consumption of AI training has also been a concern (Li et al., 2023b; Zuccon et al., 2023). Broader scholarship and policy work includes ethical (Coeckelbergh, 2021; Floridi, 2021), informational (Luccioni et al., 2024a; Stein, 2020; Pagallo et al., 2022; De Vries, 2023), and social perspectives (Bakhtiarifard et al., 2025). Numerous contributions also analyze and evaluate the various means in which AI may be used to mitigate, and adapt to, climate change (Stein, 2020; Cowls et al., 2023; Kelly, 2022; Saheb et al., 2022; Hamdan et al., 2024; Chen et al., 2023; Rolnick et al., 2022; Kumari and Pandey, 2023), as also acknowledged in the AI Act (Recitals 4 and 142). However, from a legal perspective, the problem of sustainable AI remains underexplored. Existing contributions date from before the AI Act’s final version (Pagallo et al., 2022), which differs significantly from previous proposals (see below, 5.2.) or do not engage with its provisions in detail (Hacker, 2024b). Other, complimentary work focuses uniquely on the DSA (Buri and Griffin, 2024) or data center regulation (Commins and Irion, 2025).
3. Technical Background
@@
From a technical perspective, it is important to distinguish between (pre-)training, fine-tuning, and inference. Training refers to the process of initially adjusting a model’s parameters or weights to fit the data. This process is highly compute-intensive and typically requires a significant amount of energy (Luccioni et al., 2024b, 2023; Wu et al., 2022; Kaack et al., 2022). LLM performance strongly depends on the model scale (number of parameters), which in turn requires more training data (Kaplan et al., 2020; Rae et al., 2021), and hence more resources. Such scaling can even be expected with algorithmically advanced models, such as DeepSeek R1 (Amodei, 2025).
@@
Fine-tuning aims at adjusting a pretrained model to fit more specific data in order to perform better at specific tasks. It involves a significantly smaller amount of training data and compute budget. The specific procedure for successful fine-tuning is very model-, data-, and task-specific, and rather empirical. Therefore, the computational and energy-intensity can vary significantly for this step. For example, as highlighted by Luccioni et al. (Luccioni et al., 2024b), the energy usage for fine-tuning the Bloomz-7B required 7,571 kWh compared to 51,686 kWh for the entire training process, adding another 15 % to the initial consumption. Inference, in turn, refers to the procedure of generating a prediction from a trained model. Typically, an individual inference consumes little energy compared to training; but there are many more inference than training events.
@@
Most of these training, fine-tuning and inference computations are conducted in data centers. The power usage effectiveness (PUE) metric reflects the energy efficiency of a data center. It indicates the ratio of the total energy needed by a data center, including components such as cooling, to the energy used solely by computational devices. A PUE of 1.0 would imply ideal efficiency, meaning that the data center uses only the energy necessary to power the computational devices. The average data center PUE in 2023 was 1.58 globally(Statista, 2025) and 1.6 in the EU (European Commission Joint Research Centre, 2023).
4. Regulation of Data Centers
@@
Data centers run all kinds of operations, such as cloud computing, crypto currencies and the Internet at large. Recently, AI training and inference have experienced massive growth. Broad data center regulation, therefore, indirectly governs the environmental effects of AI and constitutes the backdrop against which specific AI regulation must be viewed. Taking a look at the requirements for data centers can help to find a coherent and effective interpretation of the requirements in the AI Act as it builds on already available data and established methodology.
4.1. EU Data Collection and Reporting Obligations for Data Centers
@@
In the EU, data collection and reporting obligations for data centers were established by two recent legal acts, Art. 12 of the recast Energy Efficiency Directive EU/2023/1791 of September 13, 2023 (”EED”), and the Commission Delegated Regulation EU/2024/1364 of March 14, 2024 (”Delegated Regulation”). The new rules apply to all data centers in the EU with a power demand of the installed information technology (IT) of at least 500kW, which includes small-sized data centers. Data center operators are required to collect, make publicly available and report to a EU database information that is deemed relevant for the sustainability assessment of the data centers and the industry as a whole. The reporting is mandated on an annual basis.
@@
The required data includes energy consumption, power utilization, temperature set points, waste heat utilisation, water usage and use of renewable energy (EED, Annex VII(c)). Notably, while the EED focuses on energy and power, the reporting of water usage is a significant step forward as both energy and water consumption have raised concerns in AI settings (Li et al., 2023b; Luccioni et al., 2024b). In addition, the Water Framework Directive can be harnessed to limit the overall amount water a data center may consume, and also control for any potential loss of water quality (Hacker, 2024b).
@@
The Delegated Regulation provides specific key performance indicators and methodology. Most notable is the requirement to measure and report the energy consumption of the installed information technology. Following the standard-methodology for the calculation of PUE, [^1] The Delegated Regulation refers to the European standard CEN/CENELEC EN 50600-4-2. the energy consumption must be measured at the uninterruptible power system (UPS) or, if not existent, at the power distribution unit (PDU) or at another point specified by the data center (see Delegated Regulation, Annex II(1)(e); see also the appendix, Figure 1 as Categories 1-3.
@@
4.2. Energy Management Systems and Energy Audits
@@
The EED also requires that Member States mandate companies with an average annual energy consumption of more than 10 TJ to conduct an energy audit at least every four years and those with a consumption of more than 85 TJ to implement an energy management system including regular energy audits (Art. 11 EED). This would also apply to operators of data centers. The Directive sets up certain minimum criteria for energy audits (Annex VI EED) and refers to the relevant international or European standards (Recital 80 EED). The legal minimum criteria, however, do not dictate how energy consumption should be measured.
4.3. German 2023 Energy Efficiency Act